OpenStudy (anonymous):

Given the equation 2x+3y=4, how do you answer the following questions? (a) is the slope of the line described by this equation positive or negative? (b) As x increases in value, does y increase or decrease? (c) if x decreases by 2 units, what is the corresponding change in y?

OpenStudy (amistre64):

slope = -a/b = -2/3

OpenStudy (anonymous):

So, answer b would be decreases since it is a negative?

OpenStudy (anonymous):

So, answer b would be decreases since it is a negative?

OpenStudy (amistre64):

that is correct